# 00footnotetext: Shared first author; Shared last author

### v3

@@ Introduction
DCLM-baseline 7B 2.6T 57.1 63.7 45.4
Conclusion & limitations
We introduced the DCLM testbed and demonstrated how it leads to new state-of-the-art training sets. Our exploration of the dataset design space is only the beginning and has clear limitations. Due to compute constraints, we could only ablate design dimensions individually and could not test all approaches on larger scales. Moreover, there are many variations of DCLM-baseline that we did not explore. For instance, understanding the impact of sharded deduplication in more detail is important, and there are many more ways of training filtering models, both in terms of their architecture and training data. We also conducted most of our experiments with only one tokenizer (GPT-NeoX), and other tokenizers may perform better on multilingual tasks or math. Another limitation is that we could not sufficiently explore the run-to-run variation from different random seeds. Still, we hope that this paper is a starting point for further research on data curation that pushes the state-of-the-art beyond DCLM-baseline.
While our models trained on DCLM-baseline are competitive on common language understanding evaluations, they currently do not perform as well on code and math. We view this as a consequence of our focus on language understanding in the first version of DCLM, and not an inherent limitation of our benchmark or the DCLM-baseline training set. Indeed, prior work has shown that adding specific training data and post training methods for code and math can substantially improve performance on those domains [193, 188, 14, 90, 169]. Combining DCLM-baseline with these domain-specific training sets and extending DCLM to cover code and math are intresting avenues for future work.
There are other important performance dimensions our evaluation suite currently does not incorporate such as fairness, multilinguality, and safety. Similarly, studying toxicity or privacy filtering in the context of DCLM would be fruitful. Expanding DCLM along these dimensions is a fruitful direction for future work, and we hope that our open and accessible testbed can strengthen the foundations of data-centric research in these directions as well.
Lastly, we have only trained 7B parameter models as part of DCLM. In contrast, state-of-the-art language models are now substantially larger. While we are optimistic that our gains will also extend to larger model scales, future work still needs to test this experimentally. One possible limitation of the approach behind DCLM-baseline may be its stringent filtering ratio. After an exact global deduplication at the document level, DCLM-baseline contains approximately 2T tokens, and after removing all near-duplicates globally, about 1T tokens remain. Understanding the interaction between data quality, filtering ratio, deduplication, and multi-epoch training will be key for building larger training sets in the future.
@@ Acknowledgements.
We would like to thank Lilith Bat-Leah, Loubna Ben Allal, Samy Bengio, Mia Chiquier, Adrien Gaidon, Lizzy Grant, Tom Gunter, Awni Hannun, Jonathan Hayase, Mike Lewis, Percy Liang, Ian Magnusson, Yifan Mai, Sewon Min, David Mizrahi, Praveen Paritosh, Guilherme Penedo, Kyle Richardson, Weijia Shi, Karanjeet Singh, Joshua Susskind, Oyvind Tafjord, Carl Vondrick, and Elle Wohlmuth, for helpful feedback at various stages of the project. We would like to thank Mike Garrison and Romil Shah for help with compute and infrastructure.
This research was supported by Allen Institute for AI, Open Philanthropy, Institute for Foundations of Machine Learning (IFML), AFOSR MURI grant FA9550-22-1-0380, Israeli Science Foundation (ISF) grant no. 2486/21, Alon Fellowship, Adelis foundation, Israeli Council for Higher Education, Onassis Foundation - Scholarship ID: F ZS 056-1/2022-2023, NSF Grants AF 1901292, CNS 2148141, Tripods CCF 1934932, IFML CCF 2019844,, research gifts by Western Digital, Amazon, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco and the Stanly P. Finch Centennial Professorship in Engineering, NSF Graduate Research Fellowship. MN acknowledges funding by the Federal Ministry of Education and Research of Germany under grant no. 01IS22094B WestAI - AI Service Center West.
We gratefully acknowledge compute budget granted by Gauss Centre for Supercomputing e.V. and by the John von Neumann Institute for Computing (NIC) on the supercomputers JUWELS Booster and JURECA at Jülich Supercomputing Centre (JSC)
@@ Appendix B Additional related work
It is widely known that web-scraped text contains high quantities of boilerplate HTML, error messages, stock tickers, and other undesirable data for training language models, much of which can be detected and removed by heuristic-based systems. The exact heuristics used by each data curation method vary but can be grouped into five categories: item count, repetition count, existence, ratio, and statistics. For example, Rae et al. [128] remove any lines that contain at least two of the following stop words: the, be, to, of, and, that, have, with, defining an item count heuristic. An example of a statistic might be removing documents that have a mean line length greater than 100 characters [34].
@@
Filtering for “high-quality” data (data which was written by humans and has likely gone through an editing process [8]) is a common step for data curation pipelines. The most commonly used method for quality filtering is to train a binary classifier on data from a perceived high-quality dataset (e.g. Wikipedia) and a perceived low-quality dataset (e.g. unfiltered web text) and filter out data where the classifier assigns sufficiently low scores [59, 30, 49]. A less commonly used method is train a language model on the high-quality dataset and calculate perplexity on the data to be filtered, where high perplexity scores suggest that the data is lower quality [170, 113].
Recently, works have proposed the use of pretrained language models to identify and curate high-quality data through prompting for various dimensions of perceived quality [171, 139, 100]. Ankner et al. [12] even find that it’s possible to use a small pretrained model (125M parameters) to prune training data for models as large as 3B. MiniPile [82] demonstrated that a 1M document subset of the pile, selected by clustering and removing low-quality clusters, can lead to small LMs that maintain performance on GLUE, while significantly reducing the scale of training data. RHO-1 [92] has a similar goal to quality filtering, but rather than filtering data out of the dataset, they propose Selective Language Modeling, an objective function that selectively masks the loss of tokens that are predicted to be low quality.
@@ Appendix D Tooling
    Enrichers which always return a list of with the page as is, adding additional information to the metadata, such as detected language or number of tokens.
@@
In particular, we implement all mappers used in RefinedWeb (which includes those from Gopher as a subset) and C4 along with many new ones, and allow users to integrate custom mappers into their pipeline. Additionally, while mappers allow for document-level processing, in some cases it may also be necessary to execute corpora-level operations. For instance, a user may wish to deduplicate spans that appear in several documents. Our tooling also supports global functions that depend on all documents.
@@
Once documents have been mapped, filtered, or globally processed, we provide standardized code to tokenize and shuffle data. The output of this code is a trainable dataset artifact. For tokenization, our code uses the GPT-NeoX [24] tokenizer. Our tokenization code adopts Ray  [^2] https://github.com/ray-project/ray for cluster management and scales from a single node setups for small datasets to multiple nodes for larger ones. After tokenizing, we perform a global shuffle of our dataset.
@@
All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm. We list compute requirements for our code in Appendix Q.
@@ Appendix E DCLM-Pool
DCLM-Pool was collected by taking all 5.1M Common Crawl WARC dumps from 2008 to 2022 (inclusive) and extracting text from the html using the resiliparse framework. We opted to omit 2023 and above to prevent large amounts of language model generated text from polluting our datasets and to provide a hold out for future use. We release DCLM-Pool on HuggingFace with CC-BY-4 license. We release DCLM-Pool as a set of .jsonl files similar to Dolma-V1 and RedPajama. We provide the fields that are in the .jsonl in Table 10. The entire pool is 5.1M gzip compressed .jsonl\ files, and 340TB compressed on disk. The use of this dataset is also subject to CommonCrawl’s Terms of Use: https://commoncrawl.org/terms-of-use.
Common Crawl respects robots.txt, and thus our pool does so as well, giving content creators a mechanism to opt out of Common Crawl and DCLM-Pool. Since DCLM-Pool is a large subset of Common Crawl it will contain some PII data, however Common Crawl does honor deletion requests and periodically redacts dumps. We designed DCLM-Pool to maintain a one-to-one mapping between raw Common Crawl WARC files and DCLM-Pool .jsonl files, allowing us to update DCLM-Pool based on redactions.
We note that Common Crawl includes raw data as collected from the web without filtering. While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data. For a more complete discussion on PII and consent regarding our pools, see Appendix S.
@@ Appendix F Training details
We utilize LayerNorm [15] without bias parameters for all normalization, qk-LayerNorm [47] on queries and keys for training stability, SwiGLU [143] multilayer perceptrons (MLPs), and a depth-scaled initialization scheme following Zhang et al. [187]. Our sequence length, during pre-training is 2048. We pack multiple sequences into batches to fill the entire context, with an EOS token to split documents. We allow causal attention to attend across documents; we experimented with masking attention across documents but early experiments indicated little impact on downstream performance.
@@ Appendix G Evaluation details
  • The AGI Eval LSAT-AR dataset [194] (3-shot) tests for model knowledge in the legal domain and evaluates analytical reasoning capabilities.
@@
  • The ARC easy and ARC challenge datasets [40] (10-shot) contain four-way multiple choice questions taken from grade 3-9 science exams, where questions in the easy dataset require knowledge of basic science, and the challenge questions require some procedural reasoning.
@@
  • We use a series of 6 datasets from Big-Bench [18] (all 10-shot): (1) QA Wikidata which requires models to complete factual statements with the correct answer, (2) Dyck languages where the model needs to complete a partially balanced expression consisting of parentheses and braces, (3) Operators where the model is given some newly defined operators and asked to compute the output from some expression using those operators, (4) Repeat Copy Logic which requires the model to differentiate instructions from text-to-copy and to perform a sequence of operations, (5) CS Algorithms which requires the model to execute algorithms such as recursion and dynamic programming, and (6) Language Identification where the model is expected to identify the language of a sequence of natural language text.
@@
  • BoolQ [38] (10-shot) is a binary question answering dataset where the model is expected to answer questions about relevant passages.
@@
  • CommonsenseQA [153] (10-shot) is a 5-way multiple choice question answering dataset which evaluates the models ability to understand and apply commonsense knowledge on everyday scenarios.
@@
  • COPA [136] (0-shot) consists of causal reasoning questions where the model is given two possible outcomes to a scenario and must use commonsense to select the outcome that is more likely.
@@
  • CoQA [134] (0-shot) is a conversational question answering dataset where the model is given a passage and conversation between two participants and then expected to extract an answer from the passage to a question from one of the participants.
@@
  • HellaSwag [186] (0-shot and 10-shot) is a 4-way multiple choice commonsense reasoning dataset, where the model is required to understand implicit context and common knowledge in order to correctly select the continuation to a context.
@@
  • Jeopardy [83] (10-shot) is a dataset of questions posed in the format of the “Jeopardy!” quiz show, covering a wide variety of topics.
@@
  • LAMBADA [116] (0-shot) is a collection of narratives where a human is able to guess the final word of the narrative, but is not able to if they are only given the final sentence. To perform well on this task requires the model to attend to context from the full narrative and cannot simply rely on the local context.
@@
  • OpenBookQA [108] (0-shot) is a 4-way multiple choice question answering dataset that requires the model to use multi-step reasoning and commonsense knowledge.
@@
  • PIQA [23] (10-shot) is a binary multiple choice question answering dataset that requires the model to use physical commonsense reasoning to answer correctly.
@@
  • SQuAD [133] (10-shot) is a question answering dataset where the model is given a question and a passage containing the answer to that question.
@@
  • The Winograd Schema Challenge [89] (0-shot) is binary multiple choice pronoun resolution task where the model is given a context and asked to determine which entity a pronoun refers to, requiring the model to exhibit commonsense knowledge and contextual understanding.
@@
  • The Winogrande [140] (0-shot) dataset extends the Winograd Schema Challenge dataset by expanding the dataset to a wider variety of domains.
@@
  • We use a series of 4 additional tasks from the AGI Eval suite of datasets [194] (all 3-shot): (1) LSAT-LR and (2) LSAT-RC test for model knowledge in the legal domain and evaluate logical reasoning and reading comprehension, respectively, (3) SAT-En evaluates the model’s capabilities in English, and (4) SAT-Math evaluates the model’s capability in math using chain-of-thought prompting.
@@
  • AQuA [93] (3-shot) is a 4-way multiple choice question answering dataset that evaluates the model on algebra questions using chain-of-thought prompting.
@@
  • BBQ [117] (3-shot) is a multiple choice question answering dataset designed to detect model’s biases along nine social dimensions.
@@
  • We use a series of 9 additional datasets from Big-Bench [18] (all 10-shot): (1) Conceptual Combinations which evaluates the model’s capability to parse conceptual combinations by selecting sentences where these combinations are used correctly, (2) Conlang Translation where the model is expected to deduce a new translation from English to an obscure constructed language based on a limited number of translation examples, (3) Elementary Math QA which is a multiple choice question answering dataset of simple quantitative reasoning problems, (4) Logical Deduction which requires a model to parse, understand, and apply information about objects and relationships between objects to infer new information, (5) Misconceptions evaluates whether a model can discern popular misconceptions from truth, (6) Novel Concepts measures the models ability to creatively construct a necessary abstraction that is unlikely to have existed in training data, (7) Strange Stories measures a model’s capacity for Theory of Mind, (8) Strategy QA is a test that requires a model to answer questions requiring multi-step implicit reasoning, (9) Understanding Fables which evaluates the model’s capability to understand the moral of a short story.
@@
  • Enterprise PII classification [120] (10-shot) is a binary classification task that evaluates whether a model can detect PII (e.g. usernames, emails) within text.
@@
  • GPQA-main and GPQA-diamond [135] (5-shot) are 4-way multiple choice question answering datasets written by domain experts in biology, physics, and chemistry, which are intended to be very difficult for non-experts to answer (even with access to the web). The diamond set is a high-quality subset including only questions where two experts answer correctly, but most non-experts answer incorrectly.
@@
  • GSM8K [41] (3-shot) is a dataset of grade school math word problems that requires between 2 to 8 steps to solve, where the model uses chain-of-thought prompting.
@@
  • LogiQA [94] (10-shot) is a 4-way multiple choice question answering dataset that evaluates logical reasoning.
@@
  • Math QA [11] (10-shot) is a 5-way multiple choice question answering dataset that evaluates math word problem solving capabilities, built on top of AQuA.
@@
  • MMLU [72] (0-shot and 5-shot) is a 4-way multiple choice question answering dataset that covers 57 different domains and tasks, evaluating both world knowledge and problem solving capabilities.
@@
  • PubMedQA [78] (10-shot) is a 3-way multiple choice question answering dataset which evaluates the model’s ability to answer biomedical research questions given context from a relevant research article.
@@
  • Simple arithmetic with spaces and without spaces [110] (10-shot) are datasets consisting of simple arithmetic problems with up to 3 operations using numbers with up to 3 digits, evaluating a model’s ability to follow the correct order of operations and perform arithmetic.
@@
  • Social Interaction QA [141] (10-shot) is a binary multiple choice question answering dataset that evaluates a model’s social commonsense intelligence.
@@
  • SVAMP [119] (3-shot) is a set of challenging elementary-level math word problems that uses chain-of-thought prompting.
@@
  • Trivia QA [80] (3-shot) is an open-ended question answering dataset that evaluates the world knowledge of a model.
@@
  • The Winogender male and Winogender female datasets [138] (10-shot) are variants of the winograd schemas method that creates a minimal pair of sentences that differ only by the gender of one pronoun, designed to evaluate a model’s gender bias.
  • HumanEval [34] (0-shot) is a code completion dataset composed of hand-written problems where the model is given a function signature, and docstring and expected to correctly produce a function that will pass several unit tests.
@@
Given the multiple ways of evaluating accuracy [53], we conducted a miniature study using the LightEval evaluation framework [56]. Notably, under this framework, we are able to achieve scores above random (25%) for 0-shot MMLU for 1B models by considering the log-probabilities of entire answer passages as opposed to single letters. The 1B Hugging Face model trained on FineWeb-Edu [100] has shown to work well on this, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry. We present our findings in Figure 5.

### v4

@@ Introduction
DCLM-baseline + StarCoder + ProofPile2 7B 2.6T 57.1 63.7 45.4
Conclusion and limitations
We introduced the DCLM testbed and demonstrated how it leads to new state-of-the-art training sets. Our exploration of the dataset design space is only the beginning and has clear limitations. Due to compute constraints, we could only ablate design dimensions individually and could not test all approaches at larger scales nor train models beyond 7B parameters. We also could not sufficiently explore run-to-run variation. Moreover, there are many variations of DCLM-baseline that we did not explore, such as alternatives to sharded deduplication and using differently trained filtering models. We also conducted most of our experiments with only one tokenizer (GPT-NeoX), and other tokenizers may perform better on multilingual tasks or math. Still, we hope that this paper is a starting point for further research on data curation that pushes the state-of-the-art beyond DCLM-baseline.
While models trained on DCLM-baseline are competitive on common language understanding tasks, they currently do not perform as well on code and math. We view this as a consequence of our focus on language understanding in the first version of DCLM, and not an inherent limitation of our benchmark or dataset. Prior work has shown that adding specific training data and post training methods for code and math can substantially improve models on those domains [202, 197, 14, 96, 177]; combining DCLM-baseline with these domain-specific training sets and extending DCLM to cover code and math are interesting future directions. Other important dimensions to expand DCLM along are fairness, multilinguality, and safety. We include some analysis in Appendix S and hope that our open-source testbed can strengthen data-centric research in these directions as well.
@@ Acknowledgements.
We would like to thank Lilith Bat-Leah, Loubna Ben Allal, Samy Bengio, Mia Chiquier, Adrien Gaidon, Lizzy Grant, Tom Gunter, Awni Hannun, Tatsunori Hashimoto, Jonathan Hayase, Mike Lewis, Percy Liang, Ian Magnusson, Yifan Mai, Sewon Min, David Mizrahi, Praveen Paritosh, Guilherme Penedo, Kyle Richardson, Weijia Shi, Karanjeet Singh, Joshua Susskind, Tristan Thrush, Oyvind Tafjord, Carl Vondrick, Alexander Wettig, and Elle Wohlmuth, for helpful feedback at various stages of the project. We would also like to thank Mike Garrison and Romil Shah for help with compute and infrastructure.
This research was supported by Allen Institute for AI, Open Philanthropy, Institute for Foundations of Machine Learning (IFML), AFOSR MURI grant FA9550-22-1-0380, Israeli Science Foundation (ISF) grant no. 2486/21, Singapore AI Visiting Professorship Programme AIVP-2024-001, Alon Fellowship, Adelis foundation, Israeli Council for Higher Education, Onassis Foundation - Scholarship ID: F ZS 056-1/2022-2023, NSF Grants AF 1901292, CNS 2148141, Tripods CCF 1934932, IFML CCF 2019844,, research gifts by Western Digital, Amazon, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco and the Stanly P. Finch Centennial Professorship in Engineering, NSF Graduate Research Fellowship. MN acknowledges funding by the Federal Ministry of Education and Research of Germany under grant no. 01IS22094B WestAI - AI Service Center West.
We gratefully acknowledge compute budget granted by Gauss Centre for Supercomputing e.V. and by the John von Neumann Institute for Computing (NIC) on the supercomputers JUWELS Booster and JURECA at Jülich Supercomputing Centre (JSC).
@@ Appendix B Additional related work
It is widely known that web-scraped text contains high quantities of boilerplate HTML, error messages, stock tickers, and other undesirable data for training language models, much of which can be detected and removed by heuristic-based systems. The exact heuristics used by each data curation method vary but can be largely grouped into five categories: item count, repetition count, existence, ratio, and statistics. For example, Rae et al. [135] remove any lines that contain at least two of the following stop words: the, be, to, of, and, that, have, with, defining an item count heuristic. An example of a statistic might be removing documents that have a mean line length greater than 100 characters [37].
@@
Filtering for “high-quality” data (data which was written by humans and has likely gone through an editing process [8]) is a common step for data curation pipelines. The most commonly used method for quality filtering is to train a binary classifier on data from a perceived high-quality dataset (e.g. Wikipedia) and a perceived low-quality dataset (e.g. unfiltered web text) and filter out data where the classifier assigns sufficiently low scores [63, 31, 53]. A less commonly used method is to train a language model on the high-quality dataset and calculate perplexity on the data to be filtered, where high perplexity scores suggest that the data is lower quality [178, 119].
Recently, works have proposed the use of pretrained language models to identify and curate high-quality data through prompting for various dimensions of perceived quality [179, 146, 106]. Ankner et al. [12] even find that it’s possible to use a small pretrained model (125M parameters) to prune training data for models as large as 3B. MiniPile [88] demonstrated that a 1M document subset of the Pile selected by clustering and removing low-quality clusters, can lead to small LMs that maintain performance on GLUE, while significantly reducing the scale of training data. RHO-1 [98] has a similar goal to quality filtering, but rather than filtering data out of the dataset, they propose Selective Language Modeling, an objective function that selectively masks the loss of tokens that are predicted to be low quality.
@@ Appendix D Tooling
    Enrichers which always return a list of documents with the page content as is, adding additional information to the metadata, such as detected language or number of tokens.
@@
In particular, we implement all mappers used in RefinedWeb (which includes those from Gopher as a subset) and C4 along with many new ones. We also allow users to integrate custom mappers into their pipeline. Additionally, while mappers allow for document-level processing, in some cases it may also be necessary to execute corpora-level operations. For instance, a user may wish to deduplicate spans that appear in several documents. Our tooling also supports global functions that depend on all documents.
@@
Once documents have been mapped, filtered, or globally processed, we provide standardized code to tokenize and globally shuffle training data sequences. The output of this code is a trainable dataset artifact. For tokenization, our code uses the GPT-NeoX [24] tokenizer and is provided in two forms: (1) a Ray-based implementation  [^1] https://github.com/ray-project/ray that scales from single node setups for small datasets to multiple nodes for larger ones; (2) a Rust-based implementation that is overall more efficient but only works on single nodes.
@@
All of our results, including data processing, model training, evaluations, and plots included in this paper, are reproducible using our open-source framework and the recipes in https://datacomp.ai/dclm. We list compute requirements for our code in Appendix R. We provide a list of all 416 experiments at https://github.com/mlfoundations/dclm/blob/main/assets/DCLM_model_database.csv.
@@ Appendix E DCLM-Pool
DCLM-Pool was collected by taking all 5.1M Common Crawl WARC dumps from 2013 to 2022 (inclusive) and extracting text from the HTML using the resiliparse framework. We opted to omit 2023 and above to prevent large amounts of language model generated text from polluting our datasets and to provide a hold out for future use. The entirety of DCLM-Pool is hosted by Common Crawl [^2] https://data.commoncrawl.org/contrib/datacomp/index.html while the competition subsets can either be obtained by simply downloading specific paths from DCLM-Pool or from Huggingface (all with CC-BY-4 license). We release DCLM-Pool as a set of compressed .jsonl files similar to Dolma-V1 and RedPajama. We provide the fields that are in the .jsonl in Table 9. The entire pool is 5.1M gzip compressed .jsonl files which take up 340TB in total on disk. The use of this dataset is also subject to CommonCrawl’s Terms of Use: https://commoncrawl.org/terms-of-use.
Common Crawl respects robots.txt, and thus our pool does so as well, giving content creators a mechanism to opt out of Common Crawl and DCLM-Pool. Since DCLM-Pool is a large subset of Common Crawl it will contain some PII data. However, Common Crawl does honor deletion requests and periodically redacts dumps. We designed DCLM-Pool to maintain a one-to-one mapping between raw Common Crawl WARC files and DCLM-Pool .jsonl files, allowing us to update DCLM-Pool based on redactions.
We note that Common Crawl includes raw data as collected from the web without filtering. While some of our pools, such as DCLM-baseline, underwent some filtering of malicious URLs, none have had any special treatment for PII and sensitive content to preserve representativeness of the raw data. For a more complete discussion on PII and consent regarding our pools, see Appendix U.
@@ Appendix F Training details
We utilize LayerNorm [15] without bias parameters for all normalization, qk-LayerNorm [51] on queries and keys for training stability, SwiGLU [150] multilayer perceptrons (MLPs), and a depth-scaled initialization scheme following Zhang et al. [196]. Our sequence length, during pretraining is 2048. We pack multiple sequences into batches to fill the entire context, with an EOS token to split documents. We allow causal attention to attend across documents; we experimented with masking attention across documents but early experiments indicated little impact on downstream performance.
@@ Appendix G Evaluation details
  • The AGI Eval LSAT-AR dataset [203] (3-shot, 230 examples) tests for model knowledge in the legal domain and evaluates analytical reasoning capabilities.
@@
  • The ARC easy (2376 examples) and ARC challenge (1,172 examples) datasets [43] (10-shot) contain four-way multiple choice questions taken from grade 3-9 science exams, where questions in the easy dataset require knowledge of basic science, and the challenge questions require some procedural reasoning.
@@
  • We use a series of 6 datasets from Big-Bench [18] (all 10-shot): (1) QA Wikidata (20,321 examples) which requires models to complete factual statements with the correct answer, (2) Dyck languages (1,000 examples) where the model needs to complete a partially balanced expression consisting of parentheses and braces, (3) Operators (210 examples) where the model is given some newly defined operators and asked to compute the output from some expression using those operators, (4) Repeat Copy Logic (32 examples) which requires the model to differentiate instructions from text-to-copy and to perform a sequence of operations, (5) CS Algorithms (1,320 examples) which requires the model to execute algorithms such as recursion and dynamic programming, and (6) Language Identification (10,000 examples) where the model is expected to identify the language of a sequence of natural language text.
@@
  • BoolQ [41] (10-shot, 3,270 examples) is a binary question answering dataset where the model is expected to answer questions about relevant passages.
@@
  • CommonsenseQA [161] (10-shot, 1,221 examples) is a 5-way multiple choice question answering dataset which evaluates the models ability to understand and apply commonsense knowledge on everyday scenarios.
@@
  • COPA [143] (0-shot, 100 examples) consists of causal reasoning questions where the model is given two possible outcomes to a scenario and must use commonsense to select the outcome that is more likely.
@@
  • CoQA [141] (0-shot, 6,304 examples) is a conversational question answering dataset where the model is given a passage and conversation between two participants and then expected to extract an answer from the passage to a question from one of the participants.
@@
  • HellaSwag [195] (0-shot and 10-shot, 10,042 examples) is a 4-way multiple choice commonsense reasoning dataset, where the model is required to understand implicit context and common knowledge in order to correctly select the continuation to a context.
@@
  • Jeopardy [89] (10-shot, 2,117 examples) is a dataset of questions posed in the format of the “Jeopardy!” quiz show, covering a wide variety of topics.
@@
  • LAMBADA [123] (0-shot, 5,153 examples) is a collection of narratives where a human is able to guess the final word of the narrative, but is not able to if they are only given the final sentence. To perform well on this task requires the model to attend to context from the full narrative and cannot simply rely on the local context.
@@
  • OpenBookQA [114] (0-shot, 500 examples) is a 4-way multiple choice question answering dataset that requires the model to use multi-step reasoning and commonsense knowledge.
@@
  • PIQA [23] (10-shot, 1,838 examples) is a binary multiple choice question answering dataset that requires the model to use physical commonsense reasoning to answer correctly.
@@
  • SQuAD [140] (10-shot, 10,570 examples) is a question answering dataset where the model is given a question and a passage containing the answer to that question.
@@
  • The Winograd Schema Challenge [95] (0-shot, 273 examples) is binary multiple choice pronoun resolution task where the model is given a context and asked to determine which entity a pronoun refers to, requiring the model to exhibit commonsense knowledge and contextual understanding.
@@
  • The Winogrande [147] (0-shot, 1,267 examples) dataset extends the Winograd Schema Challenge dataset by expanding the dataset to a wider variety of domains.
@@
  • We use a series of 4 additional tasks from the AGI Eval suite of datasets [203] (all 3-shot): (1) LSAT-LR (510 examples) and (2) LSAT-RC (268 examples) test for model knowledge in the legal domain and evaluate logical reasoning and reading comprehension, respectively, (3) SAT-En (206 examples) evaluates the model’s capabilities in English, and (4) SAT-Math (220 examples) evaluates the model’s capability in math using chain-of-thought prompting.
@@
  • AQuA [99] (3-shot, 245 examples) is a 4-way multiple choice question answering dataset that evaluates the model on algebra questions using chain-of-thought prompting.
@@
  • BBQ [124] (3-shot, 55,006 examples) is a multiple choice question answering dataset designed to detect model’s biases along nine social dimensions.
@@
  • We use a series of 9 additional datasets from Big-Bench [18] (all 10-shot): (1) Conceptual Combinations (103 examples) which evaluates the model’s capability to parse conceptual combinations by selecting sentences where these combinations are used correctly, (2) Conlang Translation (164 examples) where the model is expected to deduce a new translation from English to an obscure constructed language based on a limited number of translation examples, (3) Elementary Math QA (34,313 examples) which is a multiple choice question answering dataset of simple quantitative reasoning problems, (4) Logical Deduction (1,500 examples) which requires a model to parse, understand, and apply information about objects and relationships between objects to infer new information, (5) Misconceptions (219 examples) evaluates whether a model can discern popular misconceptions from truth, (6) Novel Concepts (32 examples) measures the models ability to creatively construct a necessary abstraction that is unlikely to have existed in training data, (7) Strange Stories (174 examples) measures a model’s capacity for Theory of Mind, (8) Strategy QA (2,289 examples) is a test that requires a model to answer questions requiring multi-step implicit reasoning, (9) Understanding Fables (189 examples) which evaluates the model’s capability to understand the moral of a short story.
@@
  • Enterprise PII classification [127] (10-shot, 3,395 examples) is a binary classification task that evaluates whether a model can detect PII (e.g. usernames, emails) within text.
@@
  • GPQA-main (448 examples) and GPQA-diamond (198 examples) [142] (5-shot) are 4-way multiple choice question answering datasets written by domain experts in biology, physics, and chemistry, which are intended to be very difficult for non-experts to answer (even with access to the web). The diamond set is a high-quality subset including only questions where two experts answer correctly, but most non-experts answer incorrectly.
@@
  • GSM8K [44] (3-shot, 1,319 examples) is a dataset of grade school math word problems that requires between 2 to 8 steps to solve, where the model uses chain-of-thought prompting.
@@
  • LogiQA [100] (10-shot, 651 examples) is a 4-way multiple choice question answering dataset that evaluates logical reasoning.
@@
  • Math QA [11] (10-shot, 2,983 examples) is a 5-way multiple choice question answering dataset that evaluates math word problem solving capabilities, built on top of AQuA.
@@
  • MMLU [78] (0-shot and 5-shot, 14,042 examples) is a 4-way multiple choice question answering dataset that covers 57 different domains and tasks, evaluating both world knowledge and problem solving capabilities.
@@
  • PubMedQA [84] (10-shot, 1,000 examples) is a 3-way multiple choice question answering dataset which evaluates the model’s ability to answer biomedical research questions given context from a relevant research article.
@@
  • Simple arithmetic with spaces and without spaces [116] (10-shot, 1,000 examples) are datasets consisting of simple arithmetic problems with up to 3 operations using numbers with up to 3 digits, evaluating a model’s ability to follow the correct order of operations and perform arithmetic.
@@
  • Social Interaction QA [148] (10-shot, 1,954 examples) is a binary multiple choice question answering dataset that evaluates a model’s social commonsense intelligence.
@@
  • SVAMP [126] (3-shot, 300 examples) is a set of challenging elementary-level math word problems that uses chain-of-thought prompting.
@@
  • Trivia QA [86] (3-shot, 11,313 examples) is an open-ended question answering dataset that evaluates the world knowledge of a model.
@@
  • The Winogender male and Winogender female datasets [145] (10-shot, 60 examples) are variants of the winograd schemas method that creates a minimal pair of sentences that differ only by the gender of one pronoun, designed to evaluate a model’s gender bias.
@@
Given the multiple ways of evaluating accuracy [57], we conducted a miniature study using the LightEval evaluation framework [60]. Notably, under this framework, we are able to achieve scores above random (25%) for 0-shot MMLU for 1B models by considering the log-probabilities of entire answer passages as opposed to single letters. The 1B Hugging Face model trained on FineWeb-Edu [106] has shown to work well given this evaluation protocol, so we wanted to more closely examine how LightEval evaluation scores correlate with evaluation scores from LLM Foundry. We present our findings in Figure 5.